OpenStudy (anonymous):

a five digit no is formed using the digits 1,3,5,7,9 without repeating any of them.what is the sum of all such possible five digit nos?

OpenStudy (anonymous):

how many such numbers are there?

OpenStudy (anonymous):

they will be 120

OpenStudy (anonymous):

of the 120 how many times will the 1 be in the ten thousands place?

OpenStudy (anonymous):

probably 24

OpenStudy (anonymous):

yes exactly 24

OpenStudy (anonymous):

otherwise some other number would be there more !

OpenStudy (anonymous):

similarly the 3 will be there 24 times the 5 and 7 and 9 also

OpenStudy (anonymous):

yea therefore there sum would be 24(1+3+5+7+9)

OpenStudy (anonymous):

yes, for each place

OpenStudy (anonymous):

for the ten thousands place it would be \[10,000\times 24\times (1+3+5+7+9)\]

OpenStudy (anonymous):

better known as \[10,000\times 24\times 25\] repeat for each place

OpenStudy (anonymous):

got it thnx

OpenStudy (anonymous):

yw

OpenStudy (zarkon):

use symmetry.. \[\Large\frac{97531+13579}{2}\cdot 5!\]

OpenStudy (anonymous):

ooh much slicker

OpenStudy (anonymous):

not sure i would ever have come up with adding forwards and backwards